# Region-aware Grasp Framework with Normalized Grasp Space for Efficient 6-DoF Grasping

### v2

@@ Abstract
A series of region-based methods succeed in extracting regional features and enhancing grasp detection quality. However, faced with a cluttered scene with potential collision, the definition of the grasp-relevant region stays inconsistent, and the relationship between grasps and regional spaces remains incompletely investigated. In this paper, we propose Normalized Grasp Space (NGS) from a novel region-aware viewpoint, unifying the grasp representation within a normalized regional space and benefiting the generalizability of methods. Leveraging the NGS, we find that CNNs are underestimated for 3D feature extraction and 6-DoF grasp detection in clutter scenes and build a highly efficient Region-aware Normalized Grasp Network (RNGNet). Experiments on the public benchmark show that our method achieves significant $>20\%$ performance gains while attaining a real-time inference speed of approximately 50 FPS. Real-world cluttered scene clearance experiments underscore the effectiveness of our method. Further, human-to-robot handover and dynamic object grasping experiments demonstrate the potential of our proposed method for closed-loop grasping in dynamic scenarios.
@@ 2 Related Work
Vision-based 6-DoF Grasp Detection: Research has focused on detecting 6-DoF grasps primarily relying on single-view RGBD images or point clouds. Pioneer works ten Pas et al. [10], Liang et al. [11] mainly follow a sample-evaluation strategy, generating numerous grasp proposals and selecting optimal grasps through an evaluation model. Qin et al. [2] employ PointNet++ [12] for per-point feature extraction and direct regression of grasp parameters. Recent advancements, benefiting from large-scale datasets [3, 13, 14], adopt an end-to-end strategy. Wang et al. [6] introduces point-wise graspness to represent grasp location probabilities and aggregate regional features for rotation prediction. HGGD [15] employs 2D CNNs to generate heatmaps for grasp locations and integrate 2D semantics with 3D geometric features to configure 6D grasps. Diverging from existing approaches, our method employs a region-aware scheme forcing the network to capture scene-independent regional features, enhancing efficiency and generalization capability.
@@ 4 Dataset Experiments
Table 5: Real-world clutter clearance results. : failed to grasp all the objects.
@@ 5 Realworld Experiments
In comparison with other methods, the results presented in Table 5 demonstrate the generalization capability to real-world grasping of our framework. However, it is worth noting that HGGD struggles to generate grasps for objects with very small scales, and AnyGrasp exhibits imprecise rotation predictions, resulting in unstable object grasping and frequent drops, particularly evident when objects are stacked. Both methods are prone to failure when the captured point cloud becomes more noisy or only very partial object observations are available. In contrast, RNGNet, with its greater robustness to noise originating from the depth sensor, demonstrates the ability to predict more stable grasps with precise rotations. This effectively reduces collisions and object drops, showcasing the potential of this method.
@@
6 Conclusion
We design a novel Region-aware Grasp Framework for 6-DoF grasp detection. Normalized Grasp Space (NGS), a unified space for region-based grasp detection, is proposed to obtain consistent region representation. With the representation, we further build an efficient convolution-based Region Normalized Grasp Network (RNGNet) to extract features and predict rotation heatmap. We plan to explore integrating scene-level information from the multi-view image during the process of closed-loop grasping, which is critical for precise manipulation in complex scenes or with specific instructions. More limitations are discussed in the Section C of the Appendices.
@@ Appendix B Implemetation Details
Appendix C Limitations

### v3

@@ Abstract
A series of region-based methods succeed in extracting regional features and enhancing grasp detection quality. However, faced with a cluttered scene with potential collision, the definition of the grasp-relevant region stays inconsistent. In this paper, we propose Normalized Grasp Space (NGS) from a novel region-aware viewpoint, unifying the grasp representation within a normalized regional space and benefiting the generalizability of methods. Leveraging the NGS, we find that CNNs are underestimated for 3D feature extraction and 6-DoF grasp detection in clutter scenes and build a highly efficient Region-aware Normalized Grasp Network (RNGNet). Experiments on the public benchmark show that our method achieves significant $>20\%$ performance gains while attaining a real-time inference speed of approximately 50 FPS. Real-world cluttered scene clearance experiments underscore the effectiveness of our method. Further, human-to-robot handover and dynamic object grasping experiments demonstrate the potential of our proposed method for closed-loop grasping in dynamic scenarios.
@@ 2 Related Work
Vision-based 6-DoF Grasp Detection: Research has focused on detecting 6-DoF grasps primarily relying on single-view RGBD images or point clouds. Pioneer works ten Pas et al. [10], Liang et al. [11] mainly follow a sample-evaluation strategy, generating numerous grasp proposals and selecting optimal grasps through an evaluation model. Qin et al. [2] employ PointNet++ [12] for per-point feature extraction and direct regression of grasp parameters. Recent advancements, benefiting from large-scale datasets [3, 13, 14], adopt an end-to-end strategy. Wang et al. [6] introduces point-wise graspness to represent grasp location probabilities and aggregate regional features for rotation prediction. HGGD [15] employs 2D CNNs to generate heatmaps for grasp locations and integrate 2D semantics with 3D geometric features to configure 6D grasps. Tang et al. [16] investigates scene regions for high-quality grasp detection but ignores helpful regional characteristics. Diverging from existing approaches, our method employs a region-aware scheme forcing the network to capture scene-independent regional features, enhancing efficiency and generalization capability.
@@ 4 Dataset Experiments
Table 5: Real-world clutter clearance results. Robot performs grasp detection and executes grasping for each scene until no grasp is detected or 15 attempts are tried. means failed to grasp all the objects.
@@ 5 Realworld Experiments
Compared with other methods, results presented in Table 5 demonstrate the generalization capability to real-world grasping of our framework. However, it is worth noting that HGGD struggles to generate grasps for objects with small scales, and AnyGrasp exhibits imprecise rotation predictions, resulting in unstable grasping and frequent drops, particularly evident when objects are stacked. Both methods are prone to failure when the captured point cloud becomes more noisy or only partial object observations are available. In contrast, RNGNet, with its robustness to noise originating from the depth sensor, demonstrates the ability to predict more stable grasps with precise rotations. This effectively reduces collisions and object drops, showcasing the potential of this method.
@@
6 Limitation and Conclusion
Normalized Grasp Space (NGS), a unified space for region-based grasp detection, is proposed to obtain consistent region representation. With the representation, we further build an efficient convolution-based Region Normalized Grasp Network (RNGNet) to extract features and predict rotation heatmap. Though efficient, our method, relying on single-view images, is vulnerable to noises and can be only applied with rigid bodies. We plan to explore integrating scene-level information from the multi-view image during the process of closed-loop grasping, which is critical for precise manipulation in complex scenes or with specific instructions. More limitations are discussed in the Section C and D of the Appendices.
@@ Appendix A Additional Results
Extra Real-world Results: For more detailed performance evaluation, we conduct extra real-world experiments at different noise levels to illustrate the practical benefits and robustness of our method. In the extra real-world experiments, for more challenging grasping, we introduce more novel objects (50 objects in total vs 24 used in the manuscript) in the experiments, and more objects (6 to 12 now vs 6 to 9 in the manuscript) may be randomly placed in a clutter. Each method is tested with extra 6 scenes under different settings. The other settings (grasp order, collision check, robot control, etc.) are kept the same as in HGGD [15] and the manuscript.
Noise Deviation (m) Method Attempt Success Rate Scene Clearance Rate
0 AnyGrasp 52 / 67 = 77.6 % 5 / 6 = 83.3 %
RNGNet 53 / 60 = 88.3 % 6 / 6 = 100 %
0.01 AnyGrasp 41 / 89 = 46.1 % 2 / 6 = 33.3 %
RNGNet 48 / 80 = 60.0 % 4 / 6 = 66.7 %
0.02 AnyGrasp 8 / 90 = 8.9 % 0 / 6 = 0 %
RNGNet 39 / 88 = 44.3 % 2 / 6 = 33.3 %
Table 8: Additional Real-world Clutter Clearance Experiments.
As is shown in the table above, our method achieves an impressive success rate in attempts without additional noise, efficiently clearing all scenes. In comparison, AnyGrasp exhibits lower attempt success and scene clearance rates and may struggle to clear scenes due to repeated failures on specific objects. Our method, on the other hand, offers more viable grasps across scenes with multiple sampled patches, which benefit the scene clearance process. This is also illustrated in Figure 5 of the manuscript (AnyGrasp mainly uses GSNet for one-frame grasp detection).
As for grasp detection with noisy inputs, AnyGrasp shows a significant decline in performance as noise increases, while our method performs more consistent grasp detection results in noisy environments and is capable of clearing scenes with relatively high noise levels (0.02 m). These results demonstrate the superior robustness and reliability of our method across varying noise levels. Its ability to maintain high performance in the presence of noise makes it a more reliable choice, especially in environments where noise is a factor.
Real-world Ablations: One of the purposes of the proposed Normalized Grasp Space (NGS) is to better leverage geometric features and avoid camera viewpoint variations, keeping the receptive fields invariant to real-world camera positions. Thus, we conduct more extensive testing across various camera viewpoint to ablate the proposed NGS in the manuscript. We conduct robot grasping with a variable camera viewpoint rather than a fixed camera position by changing the camera distance to the table. All other experiment settings are the same as above.
Camera Viewpoint Method Attempt Success Rate Scene Clearance Rate
close: 0.5 m (default) RNGNet w/o NGS 52 / 66 = 78.8 % 5 / 6 = 83.3 %
RNGNet 53 / 60 = 88.3 % 6 / 6 = 100 %
medium: 0.6 to 0.8 m RNGNet w/o NGS 48 / 79 = 60.8 % 4 / 6 = 66.7 %
RNGNet 53 / 62 = 85.5 % 6 / 6 = 100 %
far: 0.8 to 1.0 m RNGNet w/o NGS 18 / 90 = 20.0 % 0 / 6 = 0 %
RNGNet 53 / 65 = 81.5 % 6 / 6 = 100 %
Camera Viewpoint: camera distance to the table surface
Attempt Success Rate = Successful Attempt Number / Total Attempt Number
Scene Clearance Rate = Cleared Scene Number / Total Scene Number
Table 9: Real-world Clutter Clearance Ablations of proposed Normalized Grasp Space.
As the results show, RNGNet with NGS demonstrates superior robustness and adaptability across different camera viewpoints. The inclusion of NGS significantly enhances the method’s ability to maintain high performance, even as the camera viewpoint moves further away. Without NGS, the performance drops considerably, especially at greater distances, making the NGS component crucial for reliable and consistent performance in varying conditions.
@@
Guidance Ablation: Though RNGNet utilizes the Grasp Heatmap Model as a part of the patch extractor by default. However, we only introduce the Grasp Heatmap Model for locating graspable regions, and our model can be easily used without the Grasp Heatmap Model. To better investigate the ability of our method to detect high-quality grasp in regions, we compare the performance of our method under different settings (with or without Grasp Heatmap Model). Without the Grasp Heatmap Model, we filter the table pointcloud by z coordinates and use farthest sampling on the foreground pointcloud to acquire region centers. As is shown in the table below, even without the guidance of the Grasp Heatmap Model, our method can still perform high-quality grasp detection for a relatively slower speed (more patches are processed).
Localization Seen Similar Novel Time1 (ms)
Heatmap 75.20 66.62 32.38 17
Foreground 74.75 66.06 33.59 30
1 Same settings as in Table 1.
Table 11: Ablation Experiments of Patch Localization Algorithm. Showing APs on Realsense split.
@@ Appendix B Implemetation Details
Target-oriented Grasping: Actually, by utilizing the proposed region-aware grasp framework, we successfully decouple the grasp location problem and grasp detection problem. As we state in the closed-loop grasping section in the manuscript, with RNGNet, we are able to detect grasps in any region regardless of the patch location or the scene structure. This implies a practical downstream application potential when there is no need to generate grasps across the scene. In short, Region-aware Grasp Framework allows a more flexible patch localization for different purposes.
@@
Appendix C Discussion
Data Efficiency: To further confirm the data efficiency of our framework, we conduct extra experiments by training methods with different portions of data. For fair comparisons, we test our method without Grasp Heatmap Model (pretrained Grasp Heatmap Model from HGGD uses all the training data and may lead to higher test results), that is, using the farthest sampling in the filtered foreground pointcloud to extract local regions. As is shown in the Table 12 below, our method achieves better data efficiency than other baselines. Even with only 10% training data, our method performs a competitive grasp detection performance.
Method 1/50 scenes 1/20 scenes 1/10 scenes Full scenes
GSNet 8.64 30.65 36.15 47.81
AnyGrasp 7 - 35 49.01
RNGNet w/o Heatmap 25.60 41.61 47.43 58.13
“1/X scenes” means annotations from X scenes are used out of total 100 training scenes.
Approximate results from Fig. 15 (Evaluation results on the GraspNet-1Billion test set when training with different portions of real-world data) in the AnyGrasp paper[9].
Table 12: Experiments of Method Data Efficiency. Showing APs on all Realsense test split with different portions of training data.
Equivariant Neural Networks: The proposed RNGNet is not inherently equivariant. We have incorporated rotation augmentation techniques to reduce the burden on the network to learn equivariance from scratch, thereby improving generalization and reducing the amount of training data required. Experiments above also show the efficacy of our proposed Region-aware Grasp Framework for offering better sample efficiency than former methods. In the future, to further improve the performance and efficiency of our approach, we may use equivariant networks in our framework.
Introduction of Scene Priors: Actually, our approach is designed to be easily integrated with widely adopted preprocessing modules to enhance results and extend functionality. Such shape completion and instance segmentations can serve as the data processing module for the input of our Region Normalized Grasp Network. Shape completion enhances object surface information, contributing to collision-free and stable grasp detection, while instance segmentation can provide valuable object pose information for more precise grasp detection. In practice, we have also explored integrating SAM (Segment Anything Model) [38] as the instance segmentation model and Grounding Dino [39] as the object detection model, enabling our method to address target-oriented and instruction-guided grasping tasks seamlessly.
Appendix D Limitations